# SpargeAttn: Accurate Sparse Attention Accelerating Any Model Inference

### v1

@@ Abstract
An efficient attention implementation is essential for large models due to its quadratic time complexity. Fortunately, attention commonly exhibits sparsity, i.e., many values in the attention map are near zero, allowing for the omission of corresponding computations. Many studies have utilized the sparse pattern to accelerate attention. However, most existing works focus on optimizing attention within specific models by exploiting certain sparse patterns of the attention map. A universal sparse attention that guarantees both the speedup and end-to-end performance of diverse models remains elusive. In this paper, we propose SpargeAttn, a universal sparse and quantized attention for any model. Our method uses a two-stage online filter: in the first stage, we rapidly and accurately predict the attention map, enabling the skip of some matrix multiplications in attention. In the second stage, we design an online softmax-aware filter that incurs no extra overhead and further skips some matrix multiplications. Experiments show that our method significantly accelerates diverse models, including language, image, and video generation, without sacrificing end-to-end metrics. The codes are available at https://github.com/thu-ml/SpargeAttn.  [^1] All experiments in this paper used SpargeAttn based on SageAttention. An updated implementation based on SageAttention2, is available at https://github.com/thu-ml/SpargeAttn. It further offers a 30% speedup over the Attention in this paper.
@@ 3.1 Sparse FlashAttention
[…]
@@ 3.2 Selective Token Compression for Sparse Prediction
Prediction. As shown in Step1 in Fig. 3, we first compute a mean cosine similarity across tokens for each block of $Q$ and $K$. Next, we compress each block into a single token by calculating a mean across tokens. Then, we compute a compressed $QK^{\top}$ using the compressed $Q$ and $K$. Finally, to prevent interference from non-self-similar blocks, i.e., the block similarity less than a hyper-parameter $\theta$, we set the corresponding values in $S$ to -$\infty$, and then obtain a compressed attention map through softmax. This algorithm can be expressed as:
@@ 3.7 HilbertCurve Permutation
[…]
@@
[…]
@@
[…]
@@ 4.3 Ablation Study and key Insights
Strategy only $M_{g}$ only $M_{pv}$ $M_{g}$ $M_{pv}$
@@
Analysis of sparsity from $M_{g}$ and $M_{pv}$. Table 6 shows the sparsity when only using $M_{g}$, only using $M_{pv}$, and using $M_{g}$$M_{pv}$ on Llama3.1 in NeedleInAHaystack task with 128K sequence length.
@@ A.1 Detailed Explain and results of permutation ablation
Table 8: The detailed description of different permutation methods.
@@ A.2 Ablation Study of Self-Similarity Judge
[…]

### v2

@@ Abstract
An efficient attention implementation is essential for large models due to its quadratic time complexity. Fortunately, attention commonly exhibits sparsity, i.e., many values in the attention map are near zero, allowing for the omission of corresponding computations. Many studies have utilized the sparse pattern to accelerate attention. However, most existing works focus on optimizing attention within specific models by exploiting certain sparse patterns of the attention map. A universal sparse attention that guarantees both the speedup and end-to-end performance of diverse models remains elusive. In this paper, we propose SpargeAttn, a universal sparse and quantized attention for any model. Our method uses a two-stage online filter: in the first stage, we rapidly and accurately predict the attention map, enabling the skip of some matrix multiplications in attention. In the second stage, we design an online softmax-aware filter that incurs no extra overhead and further skips some matrix multiplications. Experiments show that our method significantly accelerates diverse models, including language, image, and video generation, without sacrificing end-to-end metrics. The codes are available at https://github.com/thu-ml/SpargeAttn.  [^1] All experiments using SpargeAttn is based on SageAttention. An updated implementation based on SageAttention2, is available at https://github.com/thu-ml/SpargeAttn. It further offers a 30% speedup over the Attention in this paper.
@@ 3.1 Sparse FlashAttention
where $m_{ij}$ and $l_{ij}$ are $b_{q}\times 1$ vectors, which are initialized to $\infty$- ∞ and 00 respectively. The $\tilde{\sigma}()$ is an operator similar to softmax.: $m_{ij}=\max\{m_{i,j-1},\mathrm{rowmax}(S_{ij})\},~{}\widetilde{P}_{i,j}=\exp(S_{ij}-m_{ij})$. Finally, the output $O_{i}$ can be computed by $O_{i}=\mathrm{diag}(l_{ij})^{-1}O_{ij}$.
@@ 3.2 Selective Token Compression for Sparse Prediction
Prediction. As shown in Step1 in Fig. 3, we first compute a mean cosine similarity across tokens for each block of $Q$ and $K$. Next, we compress each block into a single token by calculating a mean across tokens. Then, we compute a compressed $QK^{\top}$ using the compressed $Q$ and $K$. Finally, to prevent interference from non-self-similar blocks, i.e., the block similarity less than a hyper-parameter $\theta$, we set the corresponding values in $S$ to $\infty$- ∞, and then obtain a compressed attention map through softmax. This algorithm can be expressed as:
@@ 3.7 HilbertCurve Permutation
Model (seq_len) Attention (Sparsity) Speed (TOPS)$\uparrow$ WikiText (Ppl.) $\downarrow$ Longbench $\uparrow$ InfiniteBench $\uparrow$ NIAH $\uparrow$
@@
Model (seq_len) Attention (Sparsity) Speed (TOPS)$\uparrow$ CLIPSIM $\uparrow$ CLIP-T $\uparrow$ VQA-a $\uparrow$ VQA-t $\uparrow$ FScore $\uparrow$
@@
        Model (seq_len)         Attention (Sparsity)        Speed (TOPS)$\uparrow$        FID $\downarrow$        CLIP $\uparrow$        IR $\uparrow$
@@ 4.3 Ablation Study and key Insights
Strategy only $M_{g}$ only $M_{pv}$ $M_{g}$ +$M_{pv}$
@@
Analysis of sparsity from $M_{g}$ and $M_{pv}$. Table 6 shows the sparsity when only using $M_{g}$, only using $M_{pv}$, and using $M_{g}$+$M_{pv}$ on Llama3.1 in NeedleInAHaystack task with 128K sequence length.
@@ A.1 Detailed Explain and results of permutation ablation
Table 8: The detailed description of different permutation methods.
@@ A.2 Ablation Study of Self-Similarity Judge
Model (seq_len) Attention (Sparsity) Speed (TOPS)$\uparrow$ NIAH $\uparrow$